OpenStudy (anonymous):

A circle is centered at (−4, −6) and has a radius of 3. Which of the following is the equation for this circle? (x + 4)^2 + (y + 6)^2 = 3 (x − 4)^2 + (y − 6)^2 = 9 (x − 4)^2 + (y − 6)^2 = 3 (x + 4)^2 + (y + 6)^2 = 9

OpenStudy (anonymous):

any help? stuck on this

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

Clarification: are those expression being squared?

OpenStudy (anonymous):

yes

OpenStudy (lstrasmann):

It would be D. (x-h)^2 + (y-k)^2 = r^2 is the equation of a circle.

OpenStudy (lstrasmann):

Assuming (h,k) is the center.

OpenStudy (anonymous):

Thank you

OpenStudy (lstrasmann):

Of course.

OpenStudy (victoriasushchik):

(h,k) is the center. r is the radius. So in your first example, (x+2)2+(y−3)2=10 We would rewrite that like this so that it fits the form I posted: (x−−2)2+(y−3)2=(10−−√)2 So now we see that h is -2, k is 3 and r is sqrt(10) That means that the center of the circle is (-2,3) and the radius is sqrt(10)